# Legend: Leveraging Representation Engineering to Annotate Safety Margin for Preference Datasets

### v1

@@ 5 Experiment
We conducted extensive experiments to evaluate the effectiveness of Legend. Given a preference dataset with Legend-annotated margins, we evaluate if Legend is more desirable to improve the performance of the reward model and the harmless alignment ability of the policy model, compared to other baselines (cf. Section 5.2). Furthermore, we comprehensively analyze the advantages of Legend and uncover the characteristics, exploring the impact of different Annotator LLMs and binning operations (cf. Section 5.3). [^3] We also conducted human studies to investigate the Legend annotation framework. Due to space limitations, we placed it in section A.3 of the supplementary materials .
@@ 5.1 Experimental Setup
Datasets. We testify the effectiveness of Legend via two benchmark datasets, i.e., the Harmless [1] and the Safe-RLHF [20]. Specifically, the Harmless dataset contains 12,254 training and 662 testing samples, while the Safe-RLHF dataset is divided into 9,000 training and 1,000 testing samples. Notably, the training splits are exclusively used to train the reward models. Our Legend is free from any training process.
@@
Baselines. In our research, we compared our model, Legend, with other established methods, to demonstrate its effectiveness [^4] While Llama2 uses a dataset enhanced with human-annotated margins for training its reward models, this dataset is not available to the public..
  • Origin. It refers to the vanilla preference dataset without margin.
@@
  • RewardEnsemble@$K$ [8]. It is the only existing method for automatically annotating margins. It involves training $K$ reward models, each individually assessing the difference in reward between response pairs. The final margin is calculated by averaging the reward differences obtained from these models. Considering the high time cost of training reward models, we consider $K=1,2,3$.
Implementation Details. In our experiments, we employ a range of reward models with varying parameter scales, including, Pythia (410M, 1.4B, 2.8B) [34], Qwen-chat (0.5B, 1.8B, 4B) [35], and Llama2-7B-chat [9]. For our Legend framework, the Annotator LLM is based on the Llama2-7B Base model, fine-tuned on the Alpaca dataset [32]. While we explore other model options in our ablation experiments (cf. Section 5.3), this model serves as the primary Annotator LLM. Finally, we assess the harmless alignment of policy models using the widely adopted pythia-6B-static-sft [36] and the best-of-n method [^5] Due to the high time cost and difficulty in convergence associated with PPO training, we do not use it for evaluating the performance of downstream alignment [37, 38]. [21] to validate the efficacy of various reward models, with $n=32,64,128,256$. As for the binning operation, we group the continuous margin values into 10 bins (cf. Section 5.3 for ablation studies). All experiments are carried out on a machine with an Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz and 4 A6000 GPUs.
Metrics. We first measure the accuracy of the trained reward model in identifying harmless responses. We then leverage the capabilities of GPT-4 [39] to compare responses generated by the policy models with different reward models and calculate the win rate [40].
5.2 Main Results on Reward Models and Harmless Alignment
This section explores the impact of Legend on reward model performance and the subsequent ability of policy models to generate harmless outputs. To achieve this, we train multiple reward models using preference data generated by different margin annotation methods, including Legend, to systematically evaluate their effectiveness. Considering the high cost of alignment evaluation of policy models, we randomly selected 100 questions from the test set of Safe-RLHF for downstream alignment evaluation. The results are presented in Table 1 and Figure 1. The detailed observations are provided below.
@@ 5.2 Main Results on Reward Models and Harmless Alignment
Table 1: The accuracy of reward models trained on datasets generated by different methods. We report the accuracy gain over Origin of each annotation method across various reward models (i.e., column Gains). Basically, incorporating margins into preference datasets enhances accuracy. Also, Legend delivers performance that rivals or even surpasses RewardEnsemble@K while significantly reducing the time cost (cf. Table 2 for time costs).
Harmless Origin 69.27 70.93 72.82 72.35 72.49 72.40 72.66 -
Safe-RLHF Origin $53.56$ $57.88$ $58.09$ $64.14$ $68.81$ $68.34$ $68.84$
Table 2: Time cost of different margin annotation methods on Safe-RLHF. We consider two types of reward models for RewardEnsemble@K since it requires multiple reward models for annotations.
Method Reward Model Time Cost
RewardEnsemble@1 Pythia-410M 24 mins
RewardEnsemble@2 Pythia-410M 2 $\times$ 24 mins
RewardEnsemble@3 Pythia-410M 3 $\times$ 24 mins
RewardEnsemble@1 Llama2-7B-chat 1 h 24 mins
RewardEnsemble@2 Llama2-7B-chat 2 $\times$ 1 h 24 mins
RewardEnsemble@3 Llama2-7B-chat 3 $\times$ 1 h 24 mins
Legend N/A 21 mins
Incorporating margins into preference datasets enhances the accuracy of reward model training. Our findings, shown in Table 1, all margin annotation methods consistently outperformed the baseline Origin model, regardless of the reward model architecture. In particular, the addition of margin annotations during training demonstrably enhances the accuracy of reward models, consistently improving performance by at least 1%. This impact is even more pronounced on the Safe-RLHF, where the Pyhia-1.4B and Pyhia-2.8B models trained with margin annotations achieve a remarkable 5% increase in accuracy. In addition, the results also indicate a positive correlation between the size of the reward model and its effectiveness. Larger models, with their increased capacity for learning, are better at discerning harmless situations and capturing the nuanced meaning expressed in the reference dataset. This suggests that larger reward models are more adept at learning the semantic differences between preferences, leading to improved performance.
Legend delivers performance that rivals or even surpasses RewardEnsemble@K while significantly reducing the time cost. On average, compared to Origin, Legend improves 1.28% and 1.63% of accuracy on Harmless and Safe-RLHF, respectively, comparable even outperforming some RewardEnsemble@K configurations. For instance, while Legend’s performance is only slightly lower than the more complex RewardEnsemble@3, it demonstrates a remarkable ability to achieve significantly better results for specific reward models, like Pyhia-410M and Pyhia-2.8B. More importantly, different from RewardEnsemble@K that relies on training extra $K$ reward models, our Legend significantly reduces training expenses, as evidenced in Table 2. In contrast, the performance of RewardEnsemble@K is directly linked to its time cost. The more reward models used, the higher the performance, but also the greater the computational burden and time cost. This inherent trade-off between effectiveness and cost hinders RewardEnsemble@K’s practical utility.
@@
5.3 Ablation Studies on Characteristics of Legend
@@
We consider the following ablation baselines of Legend to analyze its advantages and uncover its characteristics. In particular, we explore the impact of the binning operation and the Annotator LLMs. The results are summarized in Table 3. The detailed observations could be found below [^6] Due to similar conclusions and space limitations, we only present the results on the Safe-RLHF dataset here. Results on the Harmless dataset can be found in Section B of the supplementary materials..
@@
  • Legend w/o SMV. This variant skips the projection operation. Consequently, it utilizes the value of $\mathcal{V}_{i}^{H}$ from Equation 6 directly as the margin.
  • Legend w/o bin. This variant omits the binning operation.
@@
  • Legend w/ b_M. This variant aims to explore the impact of using different numbers of bins. We group $\mu_{i}$ from Eq.7 into $M$ bins with $M=3,5,7,10$. In our main experiments, the vanilla Legend employs 10 bins.
  • Legend w/ Llama2-13B Base. This variant employs Llama2-13B Base as the Annotator LLM.
@@
  • Legend w/ Llama2-7B Base. This variant employs Llama2-7B Base as the Annotator LLM, the same LLM used in the main experiments.
Table 3: Ablation study of Legend on Safe-RLHF. The accuracy gain over Origin of each annotation method is also reported. Projection and Binning operations in Legend can annotate the margin with the special semantic and reduce noise.
Origin N/A 53.56 57.88 58.09 68.84 -
@@
Precise safety margin characterization through SMV-based projection enhances the harmless alignment. Embedding distance often encompasses various semantic features, not just safety. In this case, as explained in Section 4.2, relying solely on embedding distance without SMV projection leads to an unreliable measure for safety semantics. Empirically, without SMV-based projection, the accuracy of Legend drops in most cases compared to the vanilla Legend (i.e., w/ b_10), with notable drops in accuracy on Pythia-410M (0.81%), Pythia-1.4B (3.99%), and Llama2-7B-chat (1.28%). This confirms the importance of precise safety margin characterization.
@@
Binning operation promotes the stability and effectiveness of Legend. As shown in Table 3, the binning operation significantly enhances the performance of the Legend method, resulting in at least a 1.42% increase in accuracy, compared to w/o bin. Its effectiveness lies in that the projected values used in Legend are not completely noise-free. Because it hypothesizes a perfectly linear representation that ignores potential inaccuracies when comparing similar magnitudes. Consequently, this can lead to unreliable margin annotations and introduce noise into the data. The binning operation within Legend effectively mitigates this issue by minimizing comparisons between similar-sized margins. By grouping values into bins, the method reduces the impact of noise, thereby enhancing the robustness of the margin annotations.
@@
The correlation between bin number and reward model performance demonstrates a scale-dependent relationship. The number of bins used in our annotation framework exhibits a scaling relationship with the size of the reward model. For smaller reward models, using fewer bins yields better performance (e.g., Legend with Pythia-410M and 3 bins). However, as the reward model size increases, using more bins becomes advantageous for improved performance (e.g., Legend with Llama2-7B-chat and 10 bins). This illustrates that smaller reward models have limited capacity to make subtle distinctions, resulting in coarser judgments about harmlessness. They are essentially restricted to broad assessments. Conversely, larger reward models possess the capacity to make more refined discriminations, enabling them to make more nuanced judgments about harmlessness.
Legend’s effectiveness could be hindered by the Annotator LLM’s ability to identify harmless responses. Table 3 shows that Legend consistently outperforms origin in most scenarios, regardless of whether the Annotator LLM is Llama2-7B or Llama2-13B. Except Legend w/ Llama2-13B Base on Pythia-410M, most of the results with Legend improve more than 1% accuracy. While Legend generally performs well, there’s a surprising pattern: Legend’s performance gains aren’t consistent when training on smaller reward models with the larger Llama2-13B Annotator. To understand this, we examined the margin distribution of Legend using different Annotator LLMs, and found that Llama2-13B’s distribution is more concentrated, suggesting it might be less adept at identifying harmless responses, shown in Figure 2. This is because Legend relies on the Annotator LLM to clearly distinguish between harmless and harmful content. In essence, for Legend to work effectively, the chosen LLM needs to be capable of distinguishing harmless responses with a high degree of certainty.
@@ 6 Conclusion
Understanding nuanced safety preferences is essential for developing robust and harmless LLMs that prioritize human well-being. Our research delves into the precise quantification of preference margins, revealing not just which harmless response is better, but by how much. This level of detail is critical for constructing reliable and accurate reward models that can discern subtle distinctions in safety, ensuring that LLMs can navigate complex situations with a nuanced understanding of risk. Inspired by recent breakthroughs in representation engineering, we introduce a novel, cost-effective framework for generating preference datasets enriched with margin annotations. Our method significantly reduces the manual effort required for labeling preference margins, allowing for the efficient creation of high-quality datasets. Through rigorous experimentation, we demonstrate the efficacy of our approach, advancing reward modeling and the harmless alignment ability of downstream LLMs. Moving forward, our method has broader applications beyond safety and could be used to understand other semantic features, such as fairness, truthfulness, and helpfulness. However, we need more readily available induction templates for these features to fully realize the potential of our method.
Appendix A More Details About Our Work
This section provides in-depth information about the code, data cards, datasets, and models that employed in our annotation framework. We also include additional experiments to demonstrate the framework’s effectiveness.
A.1 Our Code & Prompts
Our code is available at https://github.com/colfeng/Legend. It is under the Apache-2.0 license. We would like to express our gratitude to the following open-source projects, whose code significantly contributed to both the construction and evaluation of our annotation framework.
  • OpenRLHF https://github.com/OpenLLMAI/OpenRLHF. Building upon this project, we train and evaluate reward models, and also perform inference for the best-of-$n$ selection process. It is under the Apache-2.0 license.
  • MOSS-RLHF https://github.com/OpenLMLab/MOSS-RLHF. Utilizing this project, we calculate the win rate, based on GPT-4 evaluations, for the outcomes of the best-of-$n$. It is under the Apache-2.0 license.
–User Question–
query
–The Start of Assistant A’s Answer–
{model_answer1}
–The End of Assistant A’s Answer–
–The Start of Assistant B’s Answer–
{model_answer2}
–The End of Assistant B’s Answer–
A.2 The URL of Datasets and Models
Raw Preference Datasets. We testify the effectiveness of our annotation framework via two existing benchmark datasets, i.e., Harmless [1] and Safe-RLHF [20].
Our Annotated Datasets. The datasets we annotated is available at https://huggingface.co/datasets/ColFeng/safety-alignment-legend. It is under the cc-by-nc-4.0 license. More details can be found in Section A.4.
Table 4: The URL links and licenses of the models used in this paper.
Reward model URL License
Annotator LLM URL License
Policy model URL License
A.3 More Experiments

### v2

@@ 5 Experiment
We conducted extensive experiments to evaluate the effectiveness of Legend. Given a preference dataset with Legend-annotated margins, we evaluate if Legend is more desirable to improve the performance of the reward model and the harmless alignment ability of the policy model, compared to other baselines (cf. Section 5.2). Furthermore, we comprehensively analyze the advantages of Legend and uncover the characteristics, exploring the impact of different Annotator LLMs and binning operations (cf. Section 5.3). [^6] We also conduct human studies to investigate the Legend annotation framework and compare the distribution of harmless questions across different datasets to investigate the generalization of Legend. Due to space limitations, we place it in section A.2 (Technical Appendix) of the supplementary materials.
@@ 5.1 Experimental Setup
Datasets. We testify the effectiveness of Legend via two benchmark datasets containing various harmful and harmless responses: the Harmless (Bai et al. 2022a) and the Safe-RLHF (Dai et al. 2023)  [^7] While Llama2 uses a dataset with human-annotated margins for its reward models, the dataset is not available to the public.. Specifically, the Harmless dataset contains 12,254 training and 662 testing samples, while the Safe-RLHF dataset is divided into 9,000 training and 1,000 testing samples. Notably, the training splits are exclusively used to train the reward models used in existing annotation methods and for the final performance evaluation. Our Legend is free from any training process in annotation stage.
@@
Baselines. We compared our model, Legend, with other established methods, to demonstrate its effectiveness.
  • Origin refers to the preference dataset without margin.
@@
  • RewardEnsemble@$K$ (Wang et al. 2024a) is the only existing method for automatically annotating margins. It involves training $K$ reward models, each individually assessing the difference in reward between response pairs. The final margin is calculated by averaging the reward differences from these models. Considering the high time cost of training reward models, we consider $K=1,2,3$.
Dataset Method Pythia-410M Pythia-1.4B Pythia-2.8B Llama2-7B-chat Avg. Gains Annotation Time Cost
Harmless Origin 69.27 70.93 72.82 72.66 - -
RewardEnsemble@1 $70.17_{+0.90}$ $71.64_{+0.71}$ $72.11_{-0.71}$ $75.00_{+2.34}$ $0.81_{\pm 1.25}$ 25/26/41/92min
RewardEnsemble@2 $70.78_{+1.51}$ $72.25_{+1.32}$ $72.25_{-0.57}$ $75.33_{+2.67}$ $1.23_{\pm 1.34}$ 50/52/82/184min
RewardEnsemble@3 $70.03_{+0.76}$ $\textbf{73.43}_{+2.50}$ $74.29_{+1.47}$ $\textbf{75.47}_{+2.81}$ $1.89_{\pm 0.94}$ 75/78/123/276min
Legend $\textbf{72.92}_{+3.65}$ $72.92_{+1.99}$ $\textbf{74.35}_{+1.53}$ $73.70_{+1.04}$ $2.05_{\pm 1.13}$ 23min
Safe-RLHF Origin $53.56$ $57.88$ $58.09$ $68.84$
RewardEnsemble@1 $53.69_{+0.13}$ $61.03_{+3.15}$ $60.49_{+2.40}$ $69.24_{+0.40}$ $1.52_{\pm 1.49}$ 24/28/38/84min
RewardEnsemble@2 $51.86_{-1.70}$ $59.28_{+1.40}$ $62.21_{+4.12}$ $69.71_{+0.87}$ $1.17_{\pm 2.39}$ 48/56/76/168min
RewardEnsemble@3 $52.40_{-1.16}$ $\textbf{63.09}_{+5.21}$ $62.97_{+4.88}$ $70.37_{+1.53}$ $2.62_{\pm 3.02}$ 72/84/114/252min
Legend $\textbf{53.73}_{+0.17}$ $59.63_{+1.75}$ $\textbf{63.48}_{+5.39}$ $\textbf{70.88}_{+2.04}$ $2.34_{\pm 2.19}$ 21min
@@
Table 1: The accuracy of reward models trained on datasets generated by different methods. We report the accuracy gain over Origin of each annotation method across various reward models (i.e., Avg. Gains). Legend delivers performance that rivals or even surpasses RewardEnsemble@K while significantly reducing the time cost (i.e., column Annotation Time Cost). The “A/B/C/D min” means the annotation time cost of RewardEnsemble@K with reward model Pythia-410M/Pythia-1.4B/Pythia-2.8B/Llama2-7B-chat, respectively.
Implementation Details. In our experiments, we employ a range of reward models with varying parameter scales, including, Pythia (410M, 1.4B, 2.8B) (Biderman et al. 2023), Qwen-chat (0.5B, 1.8B, 4B) (Bai et al. 2023), and Llama2-7B-chat (Touvron et al. 2023). For our Legend framework, the Annotator LLM is based on the Llama2-7B Base model, fine-tuned on the Alpaca dataset (Wang et al. 2022) [^8] This can be replaced by other LLMs capable of instruction following without safe response abilities. To simplify the differentiation process of existing LLMs and ensure the controllability of the experiments, we prepare these LLMs for our experiments.. While we explore other model options in our ablation experiments (cf. Section 5.3), this model serves as the primary Annotator LLM. We also assess the harmless alignment of policy models using the widely adopted pythia-6B-static-sft (Havrilla et al. 2023) and the best-of-n method [^9] Due to the high time cost and difficulty of convergence of PPO training, we do not use it for evaluating the performance of downstream alignment (Christiano et al. 2017; Bai et al. 2022b). (Beirami et al. 2024) to validate the efficacy of various reward models, with $n=32,64,128,256$. As for the binning operation, we group the continuous margin values into 10 bins (cf. Section 5.3 for ablation studies). All experiments are carried out on a Ubuntu 22.04.3 machine with 1T memory, an Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz and 4 A6000 GPUs.
Metrics. We first measure the accuracy of the trained reward model in identifying harmless responses. We then leverage the capabilities of GPT-4 (Achiam et al. 2023) to compare responses generated by the policy models with different reward models and calculate the win rate (Dubois et al. 2024). To demonstrate our cost-effectiveness, we also record the computational cost, the time spent on annotation methods under the same device conditions.
5.2 Main Results on Harmless Alignment
This section explores the impact of Legend on reward model performance and the subsequent ability for policy models to generate harmless outputs (the downstream alignment performance evaluation). To achieve this, for the impact of Legend on reward model performance, we evaluate multiple reward models using preference data generated by different margin annotation methods, including Legend. For downstream alignment evaluation, considering the high cost of alignment with policy models, we randomly selected 100 questions from the test set of Safe-RLHF , which is a widely adopted setup in related works (Wang et al. 2024a). The results are presented in Table 1  [^10] Using the Wilcoxon signed-rank test, we find significant differences ($p<0.05$) between Legend and Origin in both datasets, indicating Legend outperforms Origin, while no significant differences ($p>0.05$) are found between Legend and RewardEnsemble@3, suggesting better or comparable performance. Due to space limitations, we place the results of reward models of Qwen in section A.3 (Technical Appendix) of the supplementary materials. They share the similar conclusions. and Figure 1. The detailed observations are provided below.
@@ 5.2 Main Results on Harmless Alignment
Incorporating margins into preference datasets enhances the accuracy of reward model training. Our findings, shown in Table 1, all margin annotation methods consistently outperformed the baseline Origin model, regardless of the reward model architecture. In particular, the addition of margin annotations during training demonstrably enhances the accuracy of reward models, consistently improving performance by at least 1%. This impact is even more pronounced on the Safe-RLHF, where the Pyhia-1.4B and Pyhia-2.8B models trained with margin annotations achieve a remarkable 5% increase in accuracy. In addition, the results indicate a positive correlation between the size of the reward model and its effectiveness. Larger models, with the increased capacity for learning, are better at discerning harmless situations and capturing the nuanced meaning expressed in the reference dataset. This suggests that larger reward models are more adept at learning the semantic differences between preferences for improving performance.
Legend delivers performance that rivals or even surpasses RewardEnsemble@K while significantly reducing the computational cost. On average, compared to Origin, Legend improves 2.05% and 2.34% of accuracy on Harmless and Safe-RLHF, respectively, comparable even outperforming some RewardEnsemble@K configurations. For instance, Legend’s performance demonstrates a remarkable ability to achieve significantly better results for specific reward models, like Pyhia-410M and Pyhia-2.8B. More importantly, different from RewardEnsemble@K that relies on training extra $K$ reward models, our Legend significantly reduces training expenses, as evidenced in Table 1. Legend’s time cost is fixed, consisting of the time taken to construct the SMV plus the time to inference and annotate the data. In contrast, RewardEnsemble@K’s time cost is determined by the training of multiple additional reward models for annotation and the inference of annotations on the data. Therefore, the performance of RewardEnsemble@K is directly linked to its time cost. The more reward models used for margin annotation, the higher the performance, but also the greater the computational burden and time cost. This inherent trade-off between effectiveness and cost hinders RewardEnsemble@K’s practical utility.
@@
5.3 In-depth Analysis on Legend
@@
We consider the following ablation baselines of Legend to analyze its advantages and uncover its characteristics. In particular, we explore the impact of the binning operation and the Annotator LLMs. The detailed observations could be found below [^11] Due to similar conclusions and space limitations, we present the results on the Safe-RLHF dataset here. More detailed results on the Safe-RLHF dataset and the Harmless dataset can be found in Section A.3 (Technical Appendix) of the supplementary materials..
@@
  • Legend w/o SMV skips the projection operation. Consequently, it utilizes the value of $\mathcal{V}_{i}^{H}$ from Equation 6 directly as the margin.
  • Legend w/o bin omits the binning operation.
@@
  • Legend w/ b_M aims to explore the impact of using different numbers of bins. We group $\mu_{i}$ from Eq.7 into $M$ bins with $M=3,5,7,10$. In our main experiments, the vanilla Legend employs 10 bins.
  • Legend w/ Llama2-13B Base employs Llama2-13B Base as the Annotator LLM.
@@
  • Legend w/ Llama2-7B Base employs Llama2-7B Base as the Annotator LLM, the same as the main experiments.
Figure 2: Results of Legend w/o SMV and Legend w/o bin.
@@
Why Legend works – Precise safety margin characterization through SMV-based projection mainly enhances the harmless alignment. Embedding distance often encompasses various semantic features, not just safety. In this case, as explained in Section 4.2, relying solely on embedding distance without SMV projection leads to an unreliable measure for safety semantics. According to Figure 2, without SMV-based projection, the accuracy of Legend drops in most cases compared to the vanilla Legend (i.e., w/ b_10), with notable drops in accuracy on Pythia-410M (0.81%), Pythia-1.4B (3.99%), and Llama2-7B-chat (1.28%). This confirms the importance of precise safety margin characterization.
@@
Is Legend stable – Binning operation used in Legend promotes the stability of Legend. As shown in Figure 2 and 3, the binning operation significantly enhances the performance of the Legend method, resulting in at least a 1.42% increase in accuracy, compared to w/o bin. Its effectiveness lies in that the projected values used in Legend are not completely noise-free. Because it hypothesizes a perfectly linear representation that ignores potential inaccuracies when comparing similar magnitudes. Consequently, this can lead to unreliable margin annotations and introduce noise into the data. The binning operation within Legend effectively mitigates this issue by minimizing comparisons between similar-sized margins. By grouping values into bins, the method reduces the impact of noise, thereby enhancing the robustness of the annotations.
Figure 3: Results of Legend w/ different binning operations.
Can it be applied with diverse reward models – Legend is flexible enough to accommodate diverse reward models of varying sizes by adjusting the number of bins. The number of bins used in Legend exhibits a scaling relationship with the size of the reward model. For smaller reward models, using fewer bins yields better performance (e.g., Legend with Pythia-410M and 3 bins in Figure 3). However, as the reward model size increases, using more bins becomes advantageous for improved performance (e.g., Legend with Llama2-7B-chat and 10 bins). This illustrates that smaller reward models have limited capacity to make subtle distinctions, resulting in coarser judgments about harmlessness. They are essentially restricted to broad assessments. Conversely, larger reward models possess the capacity to make more refined discriminations, enabling them to make nuanced judgments of harmlessness.
Method Origin Legend Legend
w/ Llama2-13B Base w/ Llama2-7B Base
Pythia-410M 53.56 52.32 53.73
Pythia-1.4B 57.88 56.39 59.63
Pythia-2.8B 58.09 58.75 63.48
Llama-7B-chat 68.84 70.49 70.88
Table 2: The results of Legend w/ different Annotator LLMs.
What is the primary bottleneck affecting the performance ceiling of LegendOur effectiveness could be hindered by the Annotator LLM’s ability to identify harmless responses. Table 2 shows that Legend consistently outperforms origin in most scenarios, regardless of whether the Annotator LLM is Llama2-7B or Llama2-13B. Except Legend w/ Llama2-13B Base on Pythia-410M, most of the results with Legend improve more than 1% accuracy. While Legend generally performs well, there’s a surprising pattern: Legend’s performance gains aren’t consistent when training on smaller reward models with the larger Llama2-13B Annotator. To understand this, we examined the margin distribution of Legend using different Annotator LLMs, and found that Llama2-13B’s distribution is more concentrated, suggesting it might be less adept at identifying harmless responses, shown in Figure 4. This is because Legend relies on the Annotator LLM to clearly distinguish between harmless and harmful content. For Legend to work effectively, the chosen LLM needs to be capable of distinguishing harmless responses with a high certainty. [^12] We also provide a heuristic method for selecting Annotator LLMs in Section A.3 (Technical Appendix) of the supplementary materials.
@@ 6 Conclusion
Understanding nuanced safety preferences is essential for developing robust and harmless LLMs that prioritize human well-being. Our research delves into the precise quantification of preference margins, revealing not just which harmless response is better, but by how much. This level of detail is critical for constructing reliable and accurate reward models that can discern subtle distinctions in safety, ensuring that LLMs can navigate complex situations with a nuanced understanding of risk. Inspired by recent breakthroughs in representation engineering, we introduce a novel, cost-effective framework for generating preference datasets enriched with margin annotations. Our method significantly reduces the manual effort required for labeling preference margins, allowing for the efficient creation of high-quality datasets. Through rigorous experimentation, we demonstrate the efficacy of our approach, advancing reward modeling and the harmless alignment ability of downstream LLMs.
Moving forward, our method has broader applications beyond safety and could be used for other semantic features, such as fairness, truthfulness, and helpfulness. However, we need more readily available induction templates for these features to fully realize the potential of our method.
Acknowledgements
This work was supported in part by the National Science and Technology Major Project (Project 2022ZD0116306); in part by the National Natural Science Foundation of China (No. 62272330); in part by the Fundamental Research Funds for the Central Universities (No. YJ202219); in part by the Science Fund for Creative Research Groups of Sichuan Province Natural Science Foundation (No. 2024NSFTD0035); in part by the National Major Scientific Instruments and Equipments Development Project of Natural Science Foundation of China under Grant (No. 62427820).
Appendix A Technical Appendix
This Technical Appendix provides more detailed results and additional experiments to demonstrate the framework’s effectiveness. We also include limitations and future works of our work.
A.1 The Visualization of the Margin Vectors of the Paired Responses
We visualize the representations of paired harmful and harmless answers (combine with the question) from the AdvBench dataset, generated by different annotator LLMs (Llama2-7B Base and Llama2-13B Base), using PCA dimensionality reduction. Figure 5 and 6 show these representations, with gray arrows (sampled from the whole dataset) indicating the relationships between pairs.
These further validates the effectiveness of our SMV. The results in the figures demonstrates that harmless and harmful answers are clearly separable into two distinct regions, and the connecting lines between pairs exhibit a consistent directional trend. This aligns with prior literature (Li et al. 2024; Qian et al. 2024), indicating that LLMs can differentiate the degree of safety between the answers and that the consistent directional trend provides an empirical basis for constructing the SMV.
Figure 5: The Visualization of the margin vectors of the paired responses on the AdvBench dataset with Llama2-7B Base annotator LLM.
Figure 6: The Visualization of the margin vectors of the paired responses on the AdvBench dataset with Llama2-13B Base annotator LLM.
A.2 More Experiments